OpenStudy (pig_cat223):

HELP PLZ!!!!! im taking a quiz and dont know how to do it... solve this system of linear equations. seperate the x- and y-values with a comma. -4x+2y=12 -13x+3y=46

OpenStudy (greatlife44):

Now so far take a look at this, we can solve this equation by substitution. say we take the top equation: we can solve for y 2y = 12+4x divide both sides by (2) Now we've got an expression for y. y = 6 + 2x

OpenStudy (greatlife44):

@pig_cat223

OpenStudy (pig_cat223):

im here.... im just super lost and confused.... :/

OpenStudy (greatlife44):

@pig_cat223 try to solve for y in this equation: follow through what I did -4x+2y = 12

OpenStudy (marcelie):

plug in the y into the 2nd equation

OpenStudy (marcelie):

for first step ur trying to get y by itself

OpenStudy (marcelie):